OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Rearrange equation 2 so that it looks like this: y=5/2+x/2 Now substitute this into the first equation. So now you have this: 2x+3(5/2+x/2)=4 Which becomes this: 2x+15/2+3x/2=4 Multiply both sides by 2 Becomes: 4x+15+3x=8 So 7x=-7 x=-1 So x=23

OpenStudy (anonymous):

Ignore the "So x=23" I didn't erase it.